# Graph Neural Networks for O-RAN Mobility Management: A Link Prediction Approach

### v2

@@ IV-D Performance comparison
The findings demonstrate the efficacy of graph-based methodologies for link prediction in large-scale networks, even without features or temporal information, relying solely on graph structure for learning. While prior studies used GNN-based regressions models for HO estimation (e.g., [5] for radio link failure and [3] for user mobility), no prior work directly applied GNN models for link prediction in this context. Although not directly comparable, our proposal achieved an F1-score 5% higher (84%) than [5] (79%), using only graph structure information. Similarly, [3] reported 85% accuracy with nodes features processing, improvable to 90% by including user’s historical mobility data. This suggests that including temporality and features would enhance our model performance, which already demonstrates high efficiency using only the graph structure for training.
V Discussion and limitations of this work
This section discusses the main results of applying GNN-based link prediction for mobility management in O-RAN, along with the limitations of our work.
V-A GNN-link prediction and trade-offs
The choice between SEAL and VGAE involves critical trade-offs for O-RAN mobility management. While SEAL is more selective in edge identification, reducing FPs crucial for minimizing unnecessary HOs and system costs, its high computational complexity limits scalability. In contrast, VGAE is computationally more efficient and better suited for large-scale networks due to its attention mechanism, exhibiting significantly faster training and inference times. However, VGAE’s performance can decline in highly dynamic environments or with increased node density, as shown in the 20k cell dataset scenario.
Real-world wireless networks are dynamic, with new UEs and shifting data distributions. As both models are transductive, they might struggle to generalize to unseen UEs and out-of-distribution inputs, highlighting the need for inductive capabilities. These challenges highlight the need for future research into inductive Spatio-Temporal GNNs, such as TGNs or GraphMixer, which can incorporate dynamic user movement patterns and network dynamics to enhance accuracy and generalization. Including temporal information in datasets is essential to fully leverage these advanced models for more precise and timely HO decisions in real-word O-RAN deployments.
V-B Lack of available datasets
A significant limitation of this work is its reliance on a single real-world dataset, whose anonymized features and structure constrain data analysis and limit the generalization of our findings. Access to large-scale, realistic datasets with temporal information remains a critical barrier for O-RAN mobility management research. Existing publicly available datasets, such as those from testbeds (e.g., Colosseum) or simulations (e.g., ns-O-RAN), lack the necessary scale, diversity, or temporal resolution for comprehensive mobility modeling and GNN training, often featuring too few UEs or relying on approximated models. Future work will focus on training and evaluating models on datasets with temporal user mobility traces, potentially leveraging emulation tools like Keysight EXata or RICTest to benchmark against other HO solutions.
V-C Low connectivity and node density
Low connectivity and node density are inherent challenges in cellular network, where users are typically connected to a single serving cell. This sparsity significantly impacts GNN models, particularly VGAE and GAE, by hindering accurate link prediction due to insufficient neighboring information. The issue is compounded by the isolated subgraphs observed in our dataset, where users connect to individual cells without broader inter-cell connectivity. A potential solution involves incorporating higher-level nodes, such as CU or DU nodes, to increase graph density, interconnect cells, and facilitate richer message passing, thereby improving embedding generation.
V-D Negative edges creation
The process of creating negative edges is critical for autoencoder-based models, as highlighted by [14, 15]. Relying on random sampling functions can erroneously introduce positive edges as negative, or duplicate negative edges across datasets, leading to incorrect training or biased evaluation metrics. Specifically, including test edges in the training set can cause overfitting and distribution shift. To mitigate these issues, future work could leverage the bipartite nature of cellular networks during negative edge creation, which may enhance the attention mechanism for cell nodes and reduce VGAE performance degradation in large-scale environments.
@@ VI Conclusions
In this article we proposed a GNN-based link prediction solution for proactive mobility management in O-RAN. Using real-world cellular data, we compared two GNN models: an autoencoder-based model and a subgraph-based model where results show the autoencoder model offers superior speed and training efficiency, especially for larger datasets, achieving competitive predictive performance with high recall and average precision. While the subgraph model excelled in precision, the autoencoder model demonstrated better generalization and recall with larger datasets, though its performance decreased with high node density. Given the ability of the autoencoder-based model to deliver accurate next-cell predictions efficiently and its potential to address challenges in mobility management, we advocate for further exploration of this approach. Future research could refine and extend these findings, paving the way for designing innovative GNN-based HO mechanisms in future networks, marking a significant paradigm shift in mobility management.

### v3

@@ IV-D Performance comparison
The findings demonstrate the efficacy of graph-based methodologies for link prediction in different networks, even without feature information (RW dataset) or temporal processing, relying solely on graph structure for learning. While prior studies used GNN-based regressions models for HO estimation (e.g.,  [12] for radio link failure and [3] for user mobility), no prior work directly applied GNN models for link prediction in this context. Although not directly comparable, our proposal achieved an F1-score 5% higher (84%) than [12] (79%), using only graph structure information. Similarly, [3] reported 85% accuracy with nodes features processing, improvable to 90% by including user’s historical mobility data. This suggests that including temporality and features would enhance our model performance, which already demonstrates high efficiency using only the graph structure for training.
V Considerations for real-world deployment
Having demonstrated the feasibility of using GNNs for HO prediction, deploying this solution in a live O-RAN environment requires addressing several key challenges. The following considerations are crucial for evolving the proposed framework from a proof-of-concept into a robust, scalable, and adaptable real-world tool.
V-A Enhancing Model Architecture for Dynamic Networks
A primary consideration for real-world deployment is the evolution of the model to handle the dynamic and temporal nature of cellular networks. The VGAE and SEAL models are transductive. They are trained on a fixed graph and struggle to generalize to new UEs entering the network or to different network environments without retraining. Furthermore, a key challenge is modeling the continuous evolution of a UE’s connection as it moves, which constantly changes the state of its wireless channel to the serving cell. This can be represented in our framework by creating multiple, time-stamped edges between the same UE-cell pair. Thereby, the model can be fed with the data related with the edge evolution over the time, and predict the edge to a new cell. Future work should explore the adaptation of inductive models, such as Spatio-Temporal GNNs, which learn transferable patterns applicable to unseen nodes. This type of models can natively learn from the sequence of the channel changes to better understand a UE’s trajectory and make more context-aware predictions.
V-B Addressing Data-Centric Challenges
The unique characteristics of cellular network data introduce further challenges that must be managed for real-world deployment. Cellular graphs are naturally sparse, with UEs connected to the serving cell. This low graph connectivity can hinder the message-passing mechanism essential for GNNs, as nodes have limited information to aggregate from their neighbors. A potential solution is to enrich the graph structure by incorporating higher-level network nodes, such as CUs and DUs, to create a more densely connected graph that facilitates richer information flow. Another critical data-centric task is the generation of negative edges for training, as a naive random sampling approach can introduce flaws like including existing links as negative examples or duplicating samples, leading to overfitting and biased evaluation. Future implementations should use a structure-aware sampling strategy to generate more realistic negative samples. Finally, GNNs demonstrate that they can operate effectively on anonymized data by learning from the graph structure. However, GNNs present a ”black box” when it comes down to the model explainability. This limitation can be addressed by integrating specialized explainability GNN frameworks, such as GNNExplainer or GraphSHAP designed to interpret GNN predictions.
V-C Ensuring Scalability and Operational Viability
For integration of our solution into a real-world scenario, scalability is the central requirement. The solution must remain computationally efficient, operate within the available budget, and grow with network size. GNN architectures differ in cost and accuracy. Subgraph based models such as SEAL can be precise on small datasets, yet their per sample search and training cost rises quickly, which limits scale. Models such as VGAE are lighter and fit better for large graphs, although performance can drop in very dense areas. To meet operator needs, a modular approach can be used. The operator network can be partitioned into regions, for example dense urban zones with frequent HOs, suburban zones with moderate mobility, and high speed corridors with fast mobility. Each region runs a model and set of hyperparameters tailored to its conditions, which keeps inference latency and memory within target limits.
@@ VI Conclusions
In this article we proposed a GNN-based link prediction solution for proactive mobility management in O-RAN. Using real-world and synthetic cellular data, we compared two GNN models: an autoencoder-based model and a subgraph-based model, with results showing that the autoencoder model offers superior speed and training efficiency, especially for larger datasets, achieving competitive predictive performance with high recall and average precision. While the subgraph model excelled in precision, the autoencoder model demonstrated better generalization and recall with larger datasets, though its performance decreased with high node density. Given the ability of the autoencoder-based model to deliver accurate next-cell predictions efficiently and its potential to address challenges in mobility management, we advocate for further exploration of this approach. Future research could refine and extend these findings, paving the way for designing innovative GNN-based HO mechanisms in future networks, marking a significant paradigm shift in mobility management.